use the technique of derivative to find dy/dx

y=3 sin^5 x


1
Expert's answer
2021-05-07T12:02:22-0400

In this case, we use the Chain Rule.



"y=3\sin^5 x = 3(\sin x)^5"


Let "u=\sin x" so that "y=3u^5"


Taking derivative of "u" with respect to "x" , we have:


"\frac{du}{dx} = \cos x"


Similarly, the derivative of "y" with respect to "u" is given as


"\frac{dy}{du} = 15u^4"


By Chain Rule,

"\frac{dy}{dx} = \frac{dy}{du} \times \frac{du}{dx}"

Therefore, "\frac{dy}{dx} = 15(\sin x)^4 \cos x"